A block of gelatin is 60 mm long, 60mm wide and 20mm high when unstressed. A force of 255N is applied tangetially to the upper surface causing a 5mm displacement relative to the lower surface. The block is placed such that the 60 mm x 60mm dimention of the block comes on the lower and upper surface. Find the shearing stress, shearing strain, and the shear modulus.


1
Expert's answer
2021-12-07T09:56:28-0500

The shearing stress


"\sigma=\frac{F}{A}=\frac{255}{0.06\cdot0.06}=70833\ (Pa)"


The shearing strain


"\gamma =\frac{\Delta x}{l}=\frac{0.005}{0.02}=0.25"


The shear modulus


"G=\frac{F/A}{\Delta x/l}=\frac{70833}{0.25}=283332\ (Pa)"